Alden Co.’s monthly unit sales and total cost data for its operating activities of the past year follow. Management wants to use these data to predict future fixed and variable costs.

   

Month Units Sold Total Cost Month Units Sold Total Cost
1 320,000 $ 160,000 7 340,000 $ 220,000
2 160,000 100,000 8 280,000 160,000
3 280,000 220,000 9 80,000 64,000
4 200,000 100,000 10 160,000 140,000
5 300,000 230,000 11 100,000 100,000
6 200,000 120,000 12 110,000 80,000

1. Estimate both the variable costs per unit and the total monthly fixed costs using the high-low method. (Do not round intermediate calculations.)

2.  Predict future total costs when sales volume is (a) 200,000 units and (b) 300,000 units.

Predictions
Sales (units) 200,000 300,000
Total $0 $0
Answer #1

1) the variable costs per unit is $0.6 and the total monthly fixed costs is $16000 using the high-low method

using high low method

activity level is units sold

variable cost per unit=(cost of highest activity level-cost of lowest activity level)/(highest activity level-lowest activity level)

from the table given in the question

cost of highest activity level=$220000

cost of lowest activity level=$64000

highest activity level=340000 unit

lowest activity level=80000 units

variable cost per unit=($220000-$64000)/(340000-80000)=$0.6 per unit

total cost=variable cost + fixed cost

fixed cost=total cost - variable cost

for the highest activity level

total cost is $220000

variable cost=$0.6*340000=$204000

fixed cost=$220000-$204000=$16000

2)

sales (units) 200000 300000
total cost $136000 $196000

total cost=variable cost+fixed cost

given

fixed cost is $16000 and variable cost per unit is $0.6

total cost for 200000 units

total cost=$16000+(0.6*200000)=$136000

total cost for 300000 units

total cost=$16000+(0.6*300000)=$196000